# A projector–rank partition theorem for exact degrees of freedom in experimental design

### v2

@@ Abstract
Degrees of freedom ($\mathrm{df}$) allocation becomes opaque once blocking, nesting, fractionation, or unequal replication disturb the balanced one-stratum classical ANOVA. We resolve this ambiguity with a $\mathrm{df}$ partition theorem. For $N$ observations and a set $\mathcal{S}$ that indexes all randomization strata in the experiment, let $\mathbf{P}_{s}$ denote the idempotent projector that averages observations within randomization stratum $s\in\mathcal{S}$ and $\mathcal{C}_{\overline{E}}$ the factorial-contrast subspace attached to the possibly aliased effect $\overline{E}$. We prove the rank identity $N-1=\sum_{s}\sum_{\overline{E}}\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$, which simultaneously diagonalizes all strata and all effects, yielding exact integer $\mathrm{df}$ for any fixed-random mixture, block structure, or replication pattern. It results in closed-form $\mathrm{df}$ tables for unbalanced designs previously reliant on numerical approximations. We introduce aliasing measures such as the $\mathrm{df}$-retention ratio $\rho(\overline{E})$, deficiency $\delta(\overline{E})$ and the variance-inflation index $\alpha(\overline{E})$ that extend Box–Hunter’s resolution concept to multi-stratum and incomplete designs, and supply immediate diagnostics of information loss. Classical results such as Cochran’s identity when $\lvert\mathcal{S}\rvert=1$ and balanced fractions, Yates’ $\mathrm{df}$ for balanced split-plots, and resolution-$R$ when $\rho(\overline{E})=0$ emerge as corollaries. The method yields calibrated error rates, boosts power by up to 60%, and outpaces bootstrap approximations.
Keywords: Stratified ANOVA, projection matrices, alias structure analysis.
@@ 1 Introduction
Classical rank additivity, first formalized by Cochran, (1941) for a single stratum of independent errors, states that for a total number $N$ of experimental observations, the (N1)(N-1) $\mathrm{df}$ split neatly among orthogonal treatment contrasts. However, the validity of Cochran’s identity is disrupted upon the introduction of a second randomization stage or the deliberate aliasing of higher-order interactions. Empirical practice tried to cope with these departures through moment-matching approximations suggested by Satterthwaite, (1946) and Kenward and Roger, (1997) or ad hoc sums-of-squares types devised by Searle et al., (1981). However, such remedies sacrifice both mathematical exactness and interpretability—often yielding non-integer, sometimes negative, $\mathrm{df}$ in heavily unbalanced layouts. Patterson and Williams, (1976) provide construction for lattices with complete and incomplete block designs, but $\mathrm{df}$ tables are only presented for balanced cases. Verbeke and Molenberghs, (2000) discuss how approximate $\mathrm{df}$ dominate certain design of experiments, despite these shortcomings.
Cornfield and Tukey, (1956) established the use of idempotent-projector calculus for ANOVA, concentrating on one-stratum models without aliasing; Hinkelmann and Kempthorne, (2005) present more recent applications of idempotent-projectors in modern experimental settings. A parallel tradition, begun by Yates, (1935) and systematized by Bailey, (2008), advocated idempotent projectors to represent randomization strata. While elegant for balanced split-plots, this projector calculus remained a craft discipline: each novel design demanded a fresh hand-drawn Hasse diagram (see, for example, (Bailey,, 2008, Chapter 4)), and unequal replication placed the method beyond reach.
@@
Simultaneously, Box and Hunter, (1961) revolutionized $2^{k-p}$ factorial planning, but its celebrated resolution measure presupposed complete balance and a single source of error, and works only for regular, fully replicated 2–level fractions, without blocking. Pedagogical projector diagrams were presented by Bailey, (2008) to account for blocking; however, it lacks closed-form $\mathrm{df}$ under imbalance. Giesbrecht and Burns, (1985) compared exact versus Satterthwaite $F$-tests in split-plot large-sample asymptotics settings. The works of Searle et al., (1992); Harville, (1997); West et al., (2014) make advances in mixed-model effects, but all these treat $\mathrm{df}$ numerically rather than through analytic rank. The literatures on blocking, mixed models, and factorial aliasing have evolved in productive isolation. To the best of our knowledge, an exact, algebraic $\mathrm{df}$ identity that is simultaneously (i) multi-stratum, (ii) alias-aware, and (iii) robust to unequal replication has not been reported.
@@
This paper reunifies those strands through a $\mathrm{df}$ partition theorem that operates at the intersection of algebra and group representation. For a set $\mathcal{S}$ that indexes all randomization strata in the experiment, let $\mathbf{P}_{s}$ denote the $N\times N$ idempotent projector that averages observations within randomization stratum $s\in\mathcal{S}$ and let $\mathcal{C}_{\overline{E}}$ denote the factorial-contrast linear subspace of $\mathbb{R}^{N}$ attached to the effect $\overline{E}$ (an alias class comprising all factorial effects that collapse into the same contrast space once the defining relations of a fractional design are imposed). By letting the family of stratum projectors $\mathbf{P}_{s}$} act jointly with the full level-permutation group $G$, we obtain a simultaneous diagonalization in which every fixed or random effect decomposes into orthogonal blocks $\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}$, where $\operatorname{Im}\mathbf{P}_{s}=\{\mathbf{P}_{s}\mathbf{v},\mathbf{v}\in\mathbb{R}^{N}\}$ is the column space of the projector $\mathbf{P}_{s}$. Essentially, when we intersect $\operatorname{Im}\mathbf{P}_{s}$ with a factorial-contrast space $\mathcal{C}_{\overline{E}}$, we isolate the part of effect $\overline{E}$ that is visible in stratum $s\in\mathcal{S}$. The dimension of this intersection $\sum\limits_{s}\sum\limits_{\overline{E}}\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$ is exactly the number $\mathrm{df}$ that the effect retains and can be tested against the mean square residing in that stratum. Crucially, the dimension of each block—and, therefore, its $\mathrm{df}$—can be written in closed form via orbit-stabilizer counts described by (Bailey,, 2008, Section 5.6), regardless of balance or aliasing.
@@
The proposed theorem immediately produces the first exact $\mathrm{df}$ tables for the workhorses of mixed-model practice including the unbalanced split-plot, row-column, lattice, and general crossed-nested hierarchies. Where software has long fallen back on Satterthwaite or Kenward–Roger numbers, we supply integer ranks computable from small Gram matrices. These formulas both generalize and strictly contain the balanced textbook results, collapsing to them when replication is equal. Beyond enumeration, we introduce three quantitative aliasing metrics: The retention ratio $\rho(\overline{E})$ measures the proportion of an effect’s ideal $\mathrm{df}$ that survive blocking and fractionation; its complement $\delta(\overline{E})$ pinpoints the exact $\mathrm{df}$ deficiency, and the variance-inflation index $\alpha(\overline{E})\geq\rho(\overline{E})^{-1}$ bounds how severely any remaining contrasts are diluted. Taken together, they extend the Box–Hunter resolution (see Box and Hunter, (1961)) from regular two-level fractions to arbitrary level structures, stratum multiplicities, and missing data. Xu and Wu, (2001) analyze the evolution of resolution metrics for asymmetrical fractional factorial designs, but our rank-based work provides an algebraic alternative. Deng and Tang, (1999) extend Box–Hunter ideas for non-regular fractional designs and backs our claim that $\rho(\overline{E})$ generalizes resolution. Our projector–rank partition theorem including $\rho$, $\delta$ and $\alpha$ indices, therefore, fill a genuine methodological gap.
@@
A trio of classical results emerges as corollaries from the projector–rank framework. (i) Single-stratum case: If the randomization consists of only one stratum $(\lvert\mathcal{S}\rvert=1,\mathbf{P}_{s_{1}}=\mathbf{I}_{N})$, the theorem reduces to Cochran’s rank-additivity identity $N-1=\sum\limits_{\overline{E}}\dim\mathcal{C}_{\overline{E}}$, reproducing exactly the fixed-effects result of Cochran, (1941). (ii) Balanced split-plot: Assigning factor $A$ to whole plots and factor $B$ to sub-plots with equal replication ($a$ levels, $b$ whole plots per level, $c$ sub-plots per whole plot) and substituting the corresponding projectors into the theorem yields the $\mathrm{df}$ table derived by Yates, (1935): (a1)(a-1) for $A$, a(b1)a(b-1) for whole-plot error, (c1)(c-1) for $B$, (a1)(c1)(a-1)(c-1) for $AB$, and a(b1)(c1)a(b-1)(c-1) for sub-plot error. (iii) Fractional-factorial resolution: For a regular $2^{k-p}$ fraction, the df-retention ratio satisfies $\rho(\overline{E})=0$ exactly when the effect order $|E|$ is at least the length of the shortest defining word. Hence, the Box–Hunter resolution $R$ of the design is $R=\min\left\{|E|:\rho(\overline{E})=0\right\}$, extending the classical definition to a rank-based statement that remains meaningful even when replication is unequal.
@@
These theoretical advances translate into immediate practical benefits. Exact numerator and denominator $\mathrm{df}$ restore nominal type-I error control to $F$-tests in highly unbalanced agronomic and industrial trials. ANOVA diagnostic tables can flag unestimable contrasts before an experiment is run, preventing costly data-collection mistakes. Optimization heuristics can now search generator sets or blocking schemes by minimizing an integer-rank alias-loss index, avoiding computationally intensive Monte Carlo simulations. The work thus unifies disparate strands of design theory into a single algebraic framework and provides practitioners with precise $\mathrm{df}$—not approximations—for complex experiments.

### v3

@@ Abstract
In many experimental designs—split-plots, blocked or nested layouts, fractional factorials, and studies with missing or unequal replication—standard ANOVA procedures no longer tell us exactly how many independent pieces of information each effect truly carries. We provide a general degrees of freedom $(\mathrm{df})$ partition theorem that resolves this ambiguity. For $N$ observations, we show that the total information in the data (i.e., N1N-1 $\mathrm{df}$) can be split exactly across experimental effects and randomization strata by projecting the data onto each stratum and counting the $\mathrm{df}$ each effect contributes there. This yields integer $\mathrm{df}$—not approximations—for any mix of fixed and random effects, blocking structures, fractionation, or imbalance. This result yields closed-form $\mathrm{df}$ tables for unbalanced split-plot, row-column, lattice, and crossed-nested designs. We introduce practical diagnostics—the $\mathrm{df}$-retention ratio $\rho$, df deficiency $\delta$, and variance-inflation index $\alpha$—that measure exactly how many $\mathrm{df}$ an effect retains under blocking or fractionation and the resulting loss of precision, thereby extending Box–Hunter’s resolution idea to multi-stratum and incomplete designs. Classical results emerge as corollaries: Cochran’s one-stratum identity; Yates’s split-plot $\mathrm{df}$; resolution-$R$ identified when an effect retains no $\mathrm{df}$. Empirical studies on split-plot and nested designs, a blocked fractional-factorial design-selection experiment, and timing benchmarks show that our approach delivers calibrated error rates, recovers information to raise power by up to 60% without additional runs, and is orders of magnitude faster than bootstrap-based $\mathrm{df}$ approximations. The framework, therefore, offers both a unifying theory and immediately useful tools for planning and analyzing complex experiments.
Keywords: Experimental design, ANOVA, degrees of freedom, projection matrices, alias analysis.
@@ 1 Introduction
Classical rank additivity, first formalized by Cochran, (1941) for a single stratum of independent errors, states that for a total number $N$ of experimental observations, the (N1)(N-1) $\mathrm{df}$ split neatly among orthogonal treatment contrasts. However, the validity of Cochran’s identity is disrupted upon the introduction of a second randomization stage or the deliberate aliasing of higher-order interactions. Empirical practice tried to cope with these departures through moment-matching approximations suggested by Satterthwaite, (1946) and Kenward and Roger, (1997) or ad hoc sums-of-squares types devised by Searle et al., (1981). However, such remedies sacrifice both mathematical exactness and interpretability—often yielding non-integer, sometimes negative, $\mathrm{df}$ in heavily unbalanced layouts. Patterson and Williams, (1976) provide construction for lattices with complete and incomplete block designs, but $\mathrm{df}$ tables are only presented for balanced cases. Verbeke and Molenberghs, (2000) discuss how approximate $\mathrm{df}$ dominate certain design of experiments, despite these shortcomings. Cornfield and Tukey, (1956) established the use of idempotent-projector calculus for ANOVA, concentrating on one-stratum models without aliasing; Hinkelmann and Kempthorne, (2005) present more recent applications of idempotent-projectors in modern experimental settings. A parallel tradition, begun by Yates, (1935) and systematized by Bailey, (2008), advocated idempotent projectors to represent randomization strata. While elegant for balanced split-plots, this projector calculus remained a craft discipline: each novel design demanded a fresh hand-drawn Hasse diagram (see, for example, (Bailey,, 2008, Chapter 4)), and unequal replication placed the method beyond reach.
@@
In fractional factorial design, Box and Hunter, (1961) revolutionized $2^{k-p}$ factorial planning, but its celebrated resolution measure presupposed complete balance and a single source of error, and works only for regular, fully replicated 2–level fractions, without blocking. Pedagogical projector diagrams were presented by Bailey, (2008) to account for blocking; however, it lacks closed-form $\mathrm{df}$ under imbalance. Giesbrecht and Burns, (1985) compared exact versus Satterthwaite $F$-tests in split-plot large-sample asymptotics settings. The works of Searle et al., (1992); Harville, (1997); West et al., (2014) make advances in mixed-model effects, but all these treat $\mathrm{df}$ numerically rather than through analytic rank. Another avenue has largely focussed on approximations of $\mathrm{df}$ under variance heterogeneity or robustness, but does not yield on an exact algebraic partition for multi-stratum designs. For instance, Zhang, (2012) developed an approximate $\mathrm{df}$ test for heteroscedastic two-way ANOVA, while Bathke, (2004) studied when the classical $F$ with usual $\mathrm{df}$ remains valid under unequal variances and non-normality. Wang et al., (2005) used modified approximate $F$-tests to size balanced mixed-ANOVA experiments; and Kulinskaya and Dollinger, (2007) proposed weighted one-way ANOVA with improved Welch-type $\mathrm{df}$. McCloud and Parmeter, (2021) provided $\mathrm{df}$ calculations via hat-matrix arguments and explicitly contrast ANOVA versus non-ANOVA $\mathrm{df}$ again in the approximation/estimation framework. In summary, the literatures on blocking, mixed models, and factorial aliasing have evolved in productive isolation. To the best of our knowledge, an exact, algebraic $\mathrm{df}$ identity that is simultaneously (i) multi-stratum, (ii) alias-aware, and (iii) robust to unequal replication is missing in the experimental design literature.
@@
This paper reunifies those strands through a $\mathrm{df}$ partition theorem that operates at the intersection of algebra and group representation. For a set $\mathcal{S}$ that indexes all randomization strata in the experiment, let $\mathbf{P}_{s}$ denote the $N\times N$ idempotent projector that averages observations within randomization stratum $s\in\mathcal{S}$ and let $\mathcal{C}_{\overline{E}}$ denote the factorial-contrast linear subspace of $\mathbb{R}^{N}$ attached to the effect $\overline{E}$ (an alias class comprising all factorial effects that collapse into the same contrast space once the defining relations of a fractional design are imposed). By letting the family of stratum projectors $\mathbf{P}_{s}$} act jointly with the full level-permutation group $G$, we obtain a simultaneous diagonalization in which every fixed or random effect decomposes into orthogonal blocks $\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}$, where $\operatorname{Im}\mathbf{P}_{s}=\{\mathbf{P}_{s}\mathbf{v},\mathbf{v}\in\mathbb{R}^{N}\}$ is the column space of the projector $\mathbf{P}_{s}$. Essentially, when we intersect $\operatorname{Im}\mathbf{P}_{s}$ with a factorial-contrast space $\mathcal{C}_{\overline{E}}$, we isolate the part of effect $\overline{E}$ that is visible in stratum $s\in\mathcal{S}$. The dimension of this intersection $\sum_{s}\sum_{\overline{E}}\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$ is exactly the number $\mathrm{df}$ that the effect retains and can be tested against the mean square residing in that stratum. The dimension of each block—and, therefore, its $\mathrm{df}$—can be written in closed form via orbit-stabilizer counts described by (Bailey,, 2008, Section 5.6), regardless of balance or aliasing.
@@
The proposed theorem immediately produces the first exact $\mathrm{df}$ tables for the workhorses of mixed-model practice including the unbalanced split-plot, row-column, lattice, and general crossed-nested hierarchies. Where software has long fallen back on Satterthwaite or Kenward–Roger numbers, we supply integer ranks computable from small Gram matrices (see, for example, (Harville,, 1997, pp. 274)). These formulas both generalize and strictly contain the balanced textbook results, collapsing to them when replication is equal. Beyond enumeration, we introduce three quantitative aliasing metrics: The retention ratio $\rho(\overline{E})$ measures the proportion of an effect’s ideal $\mathrm{df}$ that survive blocking and fractionation; its complement $\delta(\overline{E})$ pinpoints the exact $\mathrm{df}$ deficiency, and the variance-inflation index $\alpha(\overline{E})\geq\rho(\overline{E})^{-1}$ bounds how severely any remaining contrasts are diluted. Taken together, they extend the Box–Hunter resolution (see Box and Hunter, (1961)) from regular two-level fractions to arbitrary level structures, stratum multiplicities, and missing data. Xu and Wu, (2001) analyze the evolution of resolution metrics for asymmetrical fractional factorial designs, but our rank-based work provides an algebraic alternative. Deng and Tang, (1999) extend Box–Hunter ideas for non-regular fractional designs and backs our claim that $\rho(\overline{E})$ generalizes resolution. Our projector–rank partition theorem including $\rho$, $\delta$ and $\alpha$ indices, therefore, fill a genuine methodological gap.
@@
A trio of classical results emerges as corollaries from the projector–rank framework. (i) Single-stratum case: If the randomization consists of only one stratum $(\lvert\mathcal{S}\rvert=1,\mathbf{P}_{s_{1}}=\mathbf{I}_{N})$, the theorem reduces to Cochran’s rank-additivity identity $N-1=\sum_{\overline{E}}\dim\mathcal{C}_{\overline{E}}$, reproducing exactly the fixed-effects result of Cochran, (1941). (ii) Balanced split-plot: Assigning factor $A$ to whole plots and factor $B$ to sub-plots with equal replication ($a$ levels, $b$ whole plots per level, $c$ sub-plots per whole plot) and substituting the corresponding projectors into the theorem yields the $\mathrm{df}$ table derived by Yates, (1935): (a1)(a-1) for $A$, a(b1)a(b-1) for whole-plot error, (c1)(c-1) for $B$, (a1)(c1)(a-1)(c-1) for $AB$, and a(b1)(c1)a(b-1)(c-1) for sub-plot error. (iii) Fractional-factorial resolution: For a regular $2^{k-p}$ fraction, the df-retention ratio satisfies $\rho(\overline{E})=0$ exactly when the effect order $|E|$ is at least the length of the shortest defining word. Hence, the Box–Hunter resolution $R$ of the design is $R=\min\left\{|E|:\rho(\overline{E})=0\right\}$, extending the classical definition to a rank-based statement that remains meaningful even when replication is unequal.
@@
The aforementioned theoretical advances translate into immediate practical advantages. Exact numerator and denominator $\mathrm{df}$ restore nominal type-I error control to $F$-tests in highly unbalanced agronomic and industrial trials. ANOVA diagnostic tables can flag unestimable contrasts before an experiment is run, preventing costly data-collection mistakes. Optimization heuristics can now search generator sets or blocking schemes by minimizing an integer-rank alias-loss index, avoiding computationally intensive Monte Carlo simulations. The work thus unifies disparate strands of design theory into a single algebraic framework and provides practitioners with precise $\mathrm{df}$—not approximations—for complex experiments.